I need to solve for all real roots
(x-4)^2=36
and
2(sqrt)(x+4)=3

Answer by Nate(3500)

(x - 4)^2 = 36
x - 4 = +- 6
x = 4 +- 6
x = 10 and x = -2
~~~~~~~~~~~~~~~~~~~~~
2*sqrt(x + 4) = 3
sqrt(x + 4) = 3/2
x + 4 = 9/4
x = 9/4 - 16/4
x = -7/4